OpenStudy (anonymous):

Caroline wants to give 8 equal size packets of flower seeds to here friends. She had 1/2 pound of flower seeds. How much will each packet of seeds weigh?

OpenStudy (mathmate):

In pounds, divide \(\frac{1}{2}\) lb by 8. In oz, first multiply by 16 oz/lb, then divide by 8.

OpenStudy (anonymous):

i divided 1/2 by 8 and got 16 over 1 which is just 16. but i still dont understand? what else do i do? is it 16 or 1 over 16? @mathmate

OpenStudy (mathmate):

Close! \(\frac{1}{2}\div 8 =\frac{1}{2}\times \frac{1}{8} = ? \)

OpenStudy (anonymous):

OH 16/1!! thank you

OpenStudy (mathmate):

\(\large\frac{1}{2}\div 8 =\frac{1}{2}\times \frac{1}{8} = \frac{1 \times 1}{2\times 8} = ?\)

OpenStudy (anonymous):

\[\frac{ 1 }{ 16 }\] @mathmate

OpenStudy (mathmate):

Right! \(\color{blue}{\Large\checkmark}{Way~to~go!}\)

OpenStudy (anonymous):

thank you!

OpenStudy (mathmate):

You're welcome! :)